x and y
In a table if 1 is x and y is 2
In the same table 2 is x and y 8
In the same table 3 is x and y is 9
then if 10 is the value of x what is the value of y?

Answer by ankor@dixie-net.com(22740)

x and y
In a table if 1 is x and y is 2
In the same table 2 is x and y 8
In the same table 3 is x and y is 9
then if 10 is the value of x what is the value of y?
x | y
-------
1 | 2
2 | 8
3 | 9
10 | ?
:
I think this is a quadratic equation, using the form ax^2 + bx + c = y
Find a b and c
x=1
a + b + c =2
x =2
4a + 2b + c = 8
x=3
9a + 3b + c = 9
:
Subtract the 1st equation from the 2nd
4a + 2b + c = 8
a + b + c = 2
---------------eliminates c
3a + b = 6
:
Subtract the 1st equation from the 3rd
9a + 3b + c = 9
a + b + c = 2
---------------- eliminates c again
8a + 2b = 7
multiply 3a + b = 6 by 2 and subtract from the above equation
8s + 2b = 7
6a + 2b = 12
-------------eliminates b find a
2a = -5
a = -2.5
:
Find b
3(-2.5) + b = 6
-7.5 + b = 6
b = 6 + 7.5
b = 13.5
:
Find c using the 1st equation, replacing a and b
-2.5 + 13.5 + c = 2
11 + c = 2
c = 2 - 11
c = -9
:
Our equation: y = -2.5x^2 + 13.5x - 9
x = 10
y = -2.5(10^2) + 13.5(10) - 9
y = -250 + 135 - 9
y = -124 when x = 10
:
You can confirm this for each value of x, find y